Find the x intercepts
y=x^2-3x-1
Thanks

Answer by Nate(3500)

x-int: (x,0)
y = x^2 - 3x - 1
0 = x^2 - 3x - 1
1 = x^2 - 3x
13/4 = (x - 3/2)^2
+- sqrt(13)/2 = x - 3/2
[3 +- sqrt(13)]/2 = x